OpenStudy (anonymous):

a group of age 22-29 womens height with mean of 64.3 inches and standard deviation of 2.6 inches.Find the probability that her height is more than 70.5 inches.

OpenStudy (anonymous):

\[P(X>0.5)= P\left[((X-\mu)/\sigma>((70.5-64.3)/2.6) \right]\]\[=P(Z>3.38)\]\[=1-0.9996\]\[=0.0004\] remember that n=1

OpenStudy (anonymous):

can you explain step by step how you solve the problem??